# Right Splitting, Galois Correspondence, Galois Representations and Inverse Galois Problem

### v1

@@ 2.3. Application to cases in work of Arias-de-Reyna & König
Let $n,r\in\mathbb{N}$. For a prime power $q=p^{r}$, suppose the following hold.
@@
    ${\rm GL}_{2}(\mathbb{F}_{q})$ is realizable as Galois group over $\mathbb{Q}$ of extensions $E_{1},E_{2}$ which are contained in $\bar{\mathbb{Q}}$ such that $E_{1}\cap E_{2}=K$ is a degree $n$ abelian extension of $\mathbb{Q}$.
Then $(H\times H)\rtimes{\rm GL}_{2}(\mathbb{F}_{q})/H$ (with semidirect product group law as in Thm 2.4) is realizable as Galois group over $\mathbb{Q}$ where $H$ is the unique index-$n$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ such that the quotient is abelian.
@@
We have for $l=2$, $E_{1},E_{2}$ satisfying conditions of Thm 2.4 with $G_{1}=G_{2}={\rm GL}_{2}(\mathbb{F}_{q})$, and $[G_{1}:H_{1}]=n$ and $G_{1}/H_{1}$ is abelian group and $q\equiv 1$ $mod$ $n$. Therefore $H_{1}=H_{2}=H$. Since $q\equiv 1$ $mod$ $n$ and $(n,(q-1)/n)=1$, from Corollary 2.2.1, we get the required set of representatives for $H$-cosets in ${\rm GL}_{2}(\mathbb{F}_{q})$. ∎
@@
If extensions $E_{1},E_{2}$ of $\mathbb{Q}$ contained in $\bar{\mathbb{Q}}$, are linearly disjoint over an extension $K$ of $\mathbb{Q}$, then $E_{1}\cap E_{2}=K$. We get from Remark 5.1 and Corollary 4.3 of [1], that there are two Galois extensions $L_{1},L_{2}$ over $\mathbb{Q}$ with Galois group ${\rm GL}_{2}(\mathbb{F}_{p})$ which are linearly disjoint over $\mathbb{Q}(\zeta_{p})$.
@@
From Remark 5.1 and Corollary 4.3 [1], we get $E_{1},E_{2}$ satisfying conditions of previous theorem with $n=p-1$, $K=\mathbb{Q}(\zeta_{p})$ and ${\rm GL}_{2}(\mathbb{F}_{p})/H$ is cyclic group of order $p-1$. Moreover from Corollary 2.1.3, we have $H=SL_{2}(\mathbb{F}_{p})$. ∎
@@
It is proved in Thm 1.1 of [1] that for a prime $p\geq 5$, there are infinitely many locally cyclic Galois extensions of $\mathbb{Q}$ with Galois group ${\rm GL}_{2}(\mathbb{F}_{p})$, which are pairwise linearly disjoint over $\mathbb{Q}(\sqrt{p^{*}})$ where $p^{*}=(-1)^{(p-1)/2}p$.
@@
For a prime $p\geq 5$ with $p\equiv 3$ $mod$ $4$, let $H$ be the unique index-$2$ (hence normal) subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$ then $(H\times H)\rtimes{\rm GL}_{2}(\mathbb{F}_{p})/H$ (with semidirect product group law as in above theorem) is realizable as Galois group over $\mathbb{Q}$.
@@
From Thm 1.1 [1], we get $E_{1},E_{2}$ satisfying conditions of previous theorem with $n=2$. Since $p\equiv 1$ $mod$ $2$, the conditions $p\equiv 3$ $mod$ $4$ and $(2,(p-1)/2)=1$ are equivalent. Hence we are done. ∎
@@
If $p\equiv 1$ $mod$ $4$ then $(2,(p-1)/2)=2$. Hence by Corollary 2.2.1 there is no $x\in{\rm GL}_{2}(\mathbb{F}_{p})$ such that $x\in H$ and $order(x)=2$.
@@
For a prime $p\geq 5$ with $p\equiv 3$ $mod$ $4$, let $H$ be the unique index-$2$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$. Then $H$ and $H\times H$ are realizable as Galois groups over $\mathbb{Q}\sqrt{-p}$.
@@
From previous corollary and proof of Thm 2.4, we have that $H\times H$ is realizable as Galois group over $\mathbb{Q}\sqrt{p^{*}}=\mathbb{Q}\sqrt{-p}$ ($p\equiv 3$ $mod$ $4$). Since $H$ is normal in $H\times H$ (where embedding is given by $h\mapsto(h,1)$), the quotient $H$ is also realizable as Galois group over $\mathbb{Q}\sqrt{-p}$.
Proposition 2.7.
For a prime $p\geq 5$, $(PSL_{2}(\mathbb{F}_{p})\times PSL_{2}(\mathbb{F}_{p}))\rtimes\mathbb{Z}/2\mathbb{Z}$ (with semidirect product group law as in Thm 2.4) is realizable as Galois group over $\mathbb{Q}$.
Proof.
Consider $H=\{x\in{\rm GL}_{2}(\mathbb{F}_{p})\ |\ det(x)$ is a square in $\mathbb{F}_{p}^{\times}\}$, which is the unique index-$2$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ by Corollary 2.1.1. We have $SL_{2}(\mathbb{F}_{p})\subset H\subset{\rm GL}_{2}(\mathbb{F}_{p})$. Let $h\in H$ and $det(h)=\Delta_{h}^{2}$. Then $h=\Delta_{h}I\Delta_{h}^{-1}h$ where $\Delta_{h}I\in Z({\rm GL}_{2}(\mathbb{F}_{p}))$ and $\Delta_{h}^{-1}h\in SL_{2}(\mathbb{F}_{p})$. Hence $PH\cong PSL_{2}(\mathbb{F}_{p})$ where isomorphism is given by the map sending $hZ({\rm GL}_{2}(\mathbb{F}_{p}))$ to $(\Delta_{h}^{-1}h)Z(SL_{2}(\mathbb{F}_{p}))$. Since $H$ is the unique index-$2$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$ containing $Z({\rm GL}_{2}(\mathbb{F}_{p}))$, we have that $PH\cong PSL_{2}(\mathbb{F}_{p})$ is the unique index-$2$ subgroup of $P{\rm GL}_{2}(\mathbb{F}_{p})$ by one to one correspondence between subgroups of ${\rm GL}_{2}(\mathbb{F}_{p})$ containing $Z({\rm GL}_{2}(\mathbb{F}_{p}))$ and subgroups of $P{\rm GL}_{2}(\mathbb{F}_{p})$.
We have $p>3$. Let $r\neq p-1$ be a non-square element in $\mathbb{F}_{p}$. Consider $x=\begin{pmatrix}1&r+1\\ -1&-1\end{pmatrix}\not\in H$ with $x^{2}=1\ (Z({\rm GL}_{2}(\mathbb{F}_{p})))$.
From Thm 1.1 for $P{\rm GL}_{2}(\mathbb{F}_{p})$ [1], we get $E_{1},E_{2}$ satisfying conditions of Thm 2.4 for $l=2$ with $G_{1}=G_{2}=P{\rm GL}_{2}(\mathbb{F}_{p})$, and $[G_{1}:H_{1}]=2$. Therefore $H_{1}=H_{2}=PSL_{2}(\mathbb{F}_{p})$. We observe that $\{1,xZ({\rm GL}_{2}(\mathbb{F}_{p}))\}$ is required set of representatives of $PSL_{2}(\mathbb{F}_{p})$-coset in $P{\rm GL}_{2}(\mathbb{F}_{p})$. ∎
@@ 3. Galois representations, right splitting and Inverse Galois problem
3.1. Induced Galois representations
@@
Let $G={\rm Gal}(\bar{\mathbb{Q}}/\mathbb{Q})$. Let $E$ be a number field and $\Lambda$ be a prime in its ring of integers $\mathcal{O}$. Let $E_{\Lambda}$ be the completion of $E$ with respect to $\Lambda$ and $\mathcal{O}_{\Lambda}$ be its ring of integers. Let $\mathbb{F}_{\Lambda}=\mathcal{O}_{\Lambda}/\Lambda\mathcal{O}_{\Lambda}$. $(\pi,V)$ is called a Galois Representation if
  1. (1)
    $V$ is a finite dimensional vector space over $E_{\Lambda}$ and
  2. (2)
    $\pi:G\rightarrow{\rm GL}(V)$ is a continuous homomorphism.
@@
In place of $\mathbb{Q}$, we can have a similar definition for any number field $K$ with $G={\rm Gal}(\bar{K}/K)$.
Remark 3.1.2.
Let $dim_{E_{\Lambda}}(V)=m$. Now from Prop 9.3.5 in [3], there is a basis of $V$ such that $\pi(G)\subset{\rm GL}_{m}(\mathcal{O}_{\Lambda})$ for all $g\in G$. Let us consider the maps $\pi^{\prime}:G\rightarrow{\rm GL}_{m}(\mathbb{F}_{\Lambda})$ and $\tilde{\pi}:G\rightarrow P{\rm GL}_{m}(\mathbb{F}_{\Lambda})$ obtained from $\pi$ through the quotients maps $\mathcal{O}_{\Lambda}\rightarrow\mathbb{F}_{\Lambda}$ and ${\rm GL}_{m}(\mathbb{F}_{\Lambda})\rightarrow P{\rm GL}_{m}(\mathbb{F}_{\Lambda})$.
@@
A finite group $C$ is said to be realizable as Galois group over $\mathbb{Q}$ through a Galois Representation $(\pi,V)$ if $C\cong{\rm Image}(\tilde{\pi})$.
Remark 3.2.1.
@@
Let $\sigma=\pi|_{H}:H\rightarrow{\rm GL}(W)$. Then $\sigma^{\prime}:H\rightarrow{\rm GL}_{m}(\mathbb{F}_{\Lambda})$ and $\tilde{\sigma}:H\rightarrow P{\rm GL}_{m}(\mathbb{F}_{\Lambda})$ and ${\rm Image}(\tilde{\sigma})=\tilde{\pi}(H)$. Also, $[\tilde{\pi}(G):\tilde{\pi}(H)]\leq[G:H]$.
@@
Let $\{s_{i}\}_{1\leq i\leq n}$ be a set of representatives of right cosets in $H\backslash G$ and $\{w_{j}\}_{1\leq j\leq m}$ be a basis of $W$. Then we know that $\{\phi_{s_{i},w_{j}}\}_{1\leq i\leq n,1\leq j\leq m}$ is a basis for $V$ where
@@
[…]
$\begin{split}f_{1}=\phi_{1,w_{1}},\ f_{2}=\phi_{1,w_{2}},\ \dots,\ f_{m}=\phi_{1,w_{m}},\\ f_{m+1}=\phi_{s_{2},w_{1}},\ \dots,\ f_{2m}=\phi_{s_{2},w_{m}},\dots,\\ f_{(n-1)m+1}=\phi_{s_{n},w_{1}},\ \dots,\ f_{nn}=\phi_{s_{n},w_{m}}.\end{split}$
@@
Since $\pi(G)\subset{\rm GL}_{m}(\mathcal{O}_{\Lambda})$, we conclude that $\rho(G)\subset{\rm GL}_{nm}(\mathcal{O}_{\lambda})$ and $\rho^{\prime}:G\rightarrow{\rm GL}_{nm}(\mathbb{F}_{\Lambda})$ and $\tilde{\rho}:G\rightarrow P{\rm GL}_{nm}(\mathbb{F}_{\Lambda})$ can be defined.
@@
All the following results for $\rho$ and $\pi$ in this subsection also hold for $\rho^{\prime}$ and $\pi^{\prime}$ as well as $\tilde{\rho}$ and $\tilde{\pi}$.
@@
Define a map from $\rho(H)$ to $\pi(H)$ by first block projection sending $\rho(h)$ to $\pi(h)$. This is clearly a surjective group homomorphism onto $\pi(H)$. If $\pi(h)=1$, then $\pi(s_{k}hs_{k}^{-1})=\pi(s_{k})\pi(h)\pi(s_{k})^{-1}=1$ for all $k$. Hence, this map is injective too. ∎
@@
$1\rightarrow\rho(H)\rightarrow\rho(G)\rightarrow G/H\rightarrow 1$
@@
Let $G/H$ be cyclic with representatives of $H$-cosets in $G$ of the form $\{1,s,s^{2},\dots,s^{n-1}\}$. If $g=hs^{i}\in H$ for $10\leq i\leq n-$ 1, the matrix $\rho(hs^{i})$ is given by
@@
Proof.
We compute matrices $\rho(s_{i}),\rho(s_{j})$, $\rho(s_{k})$ nd $\rho(s_{i})\rho(s_{j})$.
The $p,q$-th $m\times m$ block of $\rho(s_{i})$ is $\pi(s_{p}s_{i}s_{q}^{-1})$, where $1\leq p,q\leq n$ and $q$ is such that $s_{p}s_{i}\in s_{q}H$.
The $q,r$-th $m\times m$ block of $\rho(s_{j})$ is $\pi(s_{q}s_{j}s_{r}^{-1})$, where $1\leq q,r\leq n$ and $r$ is such that $s_{q}s_{j}\in s_{r}H$.
The $p,l$-th $m\times m$ block of $\rho(s_{k})$ is $\pi(s_{p}s_{k}s_{l}^{-1})$, where $1\leq l\leq n$ and $l$ is such that $s_{p}s_{k}\in s_{l}H$.
The $p,r$-th $m\times m$ block of $\rho(s_{i})\rho(s_{j})$ is $\pi(s_{p}s_{i}s_{q}^{-1})\pi(s_{q}s_{j}s_{r}^{-1})=\pi(s_{p}s_{i}s_{j}s_{r}^{-1})$.
Since $s_{p}s_{i}\in s_{q}H$, $s_{q}s_{j}\in s_{r}H$, we get $s_{p}s_{i}s_{j}\in s_{r}H$.
Hence $\rho(s_{i})\rho(s_{j})=\rho(s_{k})\iff(\pi(s_{p}s_{i}s_{j}s_{r}^{-1})=\pi(s_{p}s_{k}s_{l}^{-1})$ and $r=l)\iff(\pi(s_{i})\pi(s_{j})=\pi(s_{k})$ and $s_{i}s_{j}\in s_{k}H$).
@@
$1\rightarrow\rho(H)\rightarrow\rho(G)\rightarrow G/H\rightarrow 1$
is right split (that is $\rho(G)\cong\rho(H)\rtimes G/H$ for some semidirect product group law) if and only if there exists a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ with $s_{1}\in H$ such that $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(G)$. (In fact it is a subgroup with same group structure as $\{s_{i}H\}_{i}=G/H$)
@@
Let $\{r_{1},r_{2},...,r_{n}\}$ be a set of representatives of $H$-cosets in $G$. If the exact sequence is right split, let $\iota$ be splitting with $\gamma\circ\iota=id_{G/H}$. Hence $\gamma\circ\iota(r_{i}H)=r_{i}H$ for each $i$. Thus for each $i$, $\iota(r_{i}H)=\rho(h_{i}r_{i})$ for some $h_{i}\in H$. Let $s_{i}=h_{i}r_{i}$ for each $i$. Hence $\{s_{1},s_{2},...,s_{n}\}$ also forms a set of representatives for $H$-cosets in $G$ with $s_{i}H=r_{i}H$ and $\iota(s_{i}H)=\rho(s_{i})$ for each $i$. Since $\iota$ is a homomorphism, we have $\iota(s_{i}H)\iota(s_{j}H)=\iota(s_{i}s_{j}H)$ for any $i,j$. Let $s_{i}s_{j}H=s_{k}H$ for some $k$. Hence $\iota(s_{i}H)\iota(s_{j}H)=\iota(s_{k}H)$ that is $\rho(s_{i})\rho(s_{j})=\rho(s_{k})$. Hence $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(G)$.
Conversely, suppose there exists a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ with $s_{1}\in H$ and $\{\rho(s_{i})\}_{i}$ forming a multiplicatively closed subset of $\rho(G)$. Let $i\neq 1$. Then $\rho(s_{1})\rho(s_{i})=\rho(s_{k})$ for some $k$. Hence, $\rho(s_{1}s_{i})=\rho(s_{k})$. Since $s_{1}\in H$, we have $s_{1}s_{i}\in s_{i}H$. Hence by matrix calculation done above, $\rho(s_{1}s_{i})=\rho(s_{k})$ is not true unless $k=i$. Hence $\rho(s_{1})\rho(s_{i})=\rho(s_{i})$. Thus $\rho(s_{1})=1$. For any $i$, consider the set $\{\rho(s_{j})\rho(s_{i})\}_{j}$ which is a permutation of the set $\{\rho(s_{j})\}_{j}$. Hence there is a $j$ such that $\rho(s_{j})\rho(s_{i})=1$. Hence $\{\rho(s_{i})\}_{i}$ forms a subgroup of $\rho(G)$.
Now by assumption $\rho(s_{i})\rho(s_{j})=\rho(s_{l})$ for some $l$. Hence $\rho(s_{i}s_{j})=\rho(s_{l})$. Suppose $s_{i}Hs_{j}H=s_{i}s_{j}H=s_{k}H$. Then by matrix calculation done above, $\rho(s_{i}s_{j})=\rho(s_{l})$ is not true unless $l=k$. Hence $\rho(s_{i}s_{j})=\rho(s_{k})$. Hence $\{\rho(s_{i})\}_{i}$ is a group with the same group structure as $\{s_{i}H\}_{i}=G/H$.
Then we can define $\iota(s_{i}H)=\rho(s_{i})$. Hence, $\iota(s_{i}H)\iota(s_{j}H)=\rho(s_{i})\rho(s_{j})=\rho(s_{k})=\iota(s_{k}H)=\iota(s_{i}Hs_{j}H)$. Hence $\iota$ is a homomorphism. Also, $\gamma(\iota(s_{i}H))=\gamma(\rho(s_{i}))=s_{i}H$ hence $\gamma\circ\iota=id_{G/H}$. Hence the exact sequence is right split. ∎
@@
We can similarly prove Remark 2.4.1 in previous section.
@@
If there exists a set of coset representatives of $H$ in $G$ which forms a multiplicatively closed set, then the above exact sequence is right split.
@@
Let $K$ be a cyclic extension of $\mathbb{Q}$ and let $G$ and $H$ be as above. Then the above exact sequence is right split if and only if there exists a set of representatives of $H$-cosets in $G$ of the form $\{1,s,s^{2},\dots,s^{n-1}\}$ with $\rho(s)^{n}=1$. Also, if these statements are true then $order(\rho(s))=n$ and $s^{n}\in H$.

### v2

@@ 2.3. Application to cases in work of Arias-de-Reyna & König
Let $n\in\mathbb{N}$. For $q$ suppose the following hold.
@@
    The group ${\rm GL}_{2}(\mathbb{F}_{q})$ is realizable as Galois group over $\mathbb{Q}$ of extensions $E_{1},E_{2}$ which are contained in $\bar{\mathbb{Q}}$ such that $E_{1}\cap E_{2}=K$ is a degree $n$ abelian extension of $\mathbb{Q}$.
Then $(H\times H)\rtimes{\rm GL}_{2}(\mathbb{F}_{q})/H$ (with semidirect product group law as in Theorem 2.4) is realizable as Galois group over $\mathbb{Q}$ where $H$ is the unique index-$n$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ such that the quotient is abelian.
@@
We have for $l=2$, $E_{1},E_{2}$ satisfying conditions of Theorem 2.4 with $G_{1}=G_{2}={\rm GL}_{2}(\mathbb{F}_{q})$, and $G_{1}:H_{1}]=n$ and $G_{1}/H_{1}$ is abelian group and $q\equiv 1$ modmod $n$. Therefore $H_{1}=H_{2}=H$. Since $q\equiv 1$ modmod $n$ and (n,(q1)/n)=1(n,(q-1)/n)=1, from Corollary 2.2.1, we get the required set of representatives for $H$-cosets in ${\rm GL}_{2}(\mathbb{F}_{q})$.∎
@@
If extensions $E_{1},E_{2}$ of $\mathbb{Q}$ contained in $\bar{\mathbb{Q}}$, are linearly disjoint over an extension $K$ of $\mathbb{Q}$, then $E_{1}\cap E_{2}=K$. We get from [1, Remark 4.4] and [1, Corollary 4.3], that there are two Galois extensions $L_{1},L_{2}$ over $\mathbb{Q}$ with Galois group ${\rm GL}_{2}(\mathbb{F}_{p})$ which are linearly disjoint over $\mathbb{Q}(\zeta_{p})$.
@@
From [1, Remark 4.4] and [1, Corollary 4.3], we get $E_{1},E_{2}$ satisfying conditions of previous theorem with n=p1n=p-1, $K=\mathbb{Q}(\zeta_{p})$ and ${\rm GL}_{2}(\mathbb{F}_{p})/H$ is cyclic group of order p1p-1. Moreover from Corollary 2.1.2 (2), we have $H=SL_{2}(\mathbb{F}_{p})$. ∎
From previous corollary and proof of Theorem 2.4 we have the following.
Corollary 2.6.2.
For a prime $p\geq 5$, $SL_{2}(\mathbb{F}_{p})$ and $SL_{2}(\mathbb{F}_{p})\times SL_{2}(\mathbb{F}_{p})$ are realizable as Galois groups over $\mathbb{Q}(\zeta_{p})$.
@@
It is proved in [1, Theorem 1.1] that for a prime $p\geq 5$, there are infinitely many locally cyclic Galois extensions of $\mathbb{Q}$ with Galois group ${\rm GL}_{2}(\mathbb{F}_{p})$, which are pairwise linearly disjoint over $\mathbb{Q}(\sqrt{p^{*}})$ where $p^{*}=(-1)^{(p-1)/2}p$.
@@
For a prime $p\geq 5$ with $p\equiv 3$ modmod 44, let $H$ be the unique index-$2$ (hence normal) subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$. Then $(H\times H)\rtimes{\rm GL}_{2}(\mathbb{F}_{p})/H$ (with semidirect product group law as in above theorem) is realizable as Galois group over $\mathbb{Q}$.
@@
From [1, Theorem 1.1], we get $E_{1},E_{2}$ satisfying conditions of previous theorem with $n=2$. Since $p\equiv 1$ modmod 22, the conditions $p\equiv 3$ modmod 44 and (2,(p1)/2)=1(2,(p-1)/2)=1 are equivalent. Hence we are done. ∎
@@
If $p\equiv 1$ modmod 44 then (2,(p1)/2)=2(2,(p-1)/2)=2. Hence by Corollary 2.2.1 there is no $x\in{\rm GL}_{2}(\mathbb{F}_{p})$ such that $x\in H$ and $order(x)=2$.
Now for $p\equiv 3$ modmod 44 we have $p^{*}=-p$. From previous corollary and proof of Theorem 2.4 we have the following.
@@
For a prime $p\geq 5$ with $p\equiv 3$ modmod 44, let $H$ be the unique index-$2$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$. Then $H$ and $H\times H$ are realizable as Galois groups over $\mathbb{Q}\sqrt{-p}$.
Proposition 2.7.
For a prime $p\geq 5$, $(PSL_{2}(\mathbb{F}_{p})\times PSL_{2}(\mathbb{F}_{p}))\rtimes\mathbb{Z}/2\mathbb{Z}$ (with semidirect product group law as in Theorem 2.4) is realizable as Galois group over $\mathbb{Q}$.
@@
Consider $H=\{x\in{\rm GL}_{2}(\mathbb{F}_{p})\ |\ det(x)$ is a square in $\mathbb{F}_{p}^{\times}\}$, which is the unique index-$2$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{q})$ by Corollary 2.1.1. It can be seen that $H=Z({\rm GL}_{2}(\mathbb{F}_{p}))SL_{2}(\mathbb{F}_{p})$. Hence it follows that $PH\cong PSL_{2}(\mathbb{F}_{p})$. Since $H$ is the unique index-$2$ subgroup of ${\rm GL}_{2}(\mathbb{F}_{p})$ containing $Z({\rm GL}_{2}(\mathbb{F}_{p}))$, we have that $PH\cong PSL_{2}(\mathbb{F}_{p})$ is the unique index-$2$ subgroup of ${\rm PGL}_{2}(\mathbb{F}_{p})$.
We have $p\geq 5$. Let $r\neq p-1$ be a non-square element in $\mathbb{F}_{p}$. Consider $x=\begin{pmatrix}1&r+1\\ -1&-1\end{pmatrix}\not\in H$ with $x^{2}=1\ (Z({\rm GL}_{2}(\mathbb{F}_{p})))$.
From [1], we have that [1, Theorem 1.1] is also true for ${\rm PGL}_{2}(\mathbb{F}_{p})$. Thus we get $E_{1},E_{2}$ satisfying conditions of Theorem 2.4 for $l=2$ with $G_{1}=G_{2}={\rm PGL}_{2}(\mathbb{F}_{p})$, and $G_{1}:H_{1}]=2$. Therefore $H_{1}=H_{2}=PSL_{2}(\mathbb{F}_{p})$. We observe that $\{1,xZ({\rm GL}_{2}(\mathbb{F}_{p}))\}$ is required set of representatives of $PSL_{2}(\mathbb{F}_{p})$-coset in ${\rm PGL}_{2}(\mathbb{F}_{p})$. ∎
@@ 3. Galois Representations, Right Splitting and Inverse Galois Problem
3.1. Induced Galois representations and Galois groups
@@
Let $G={\rm Gal}(\bar{\mathbb{Q}}/\mathbb{Q})$. Let $E$ be a number field and $\Lambda$ be a prime in its ring of integers $\mathcal{O}$. Let $E_{\Lambda}$ be the completion of $E$ with respect to $\Lambda$ and $\mathcal{O}_{\Lambda}$ be its ring of integers. Let $\mathbb{F}_{\Lambda}=\mathcal{O}_{\Lambda}/\Lambda\mathcal{O}_{\Lambda}$.
We call $(\pi,V)$ a Galois representation if $V$ is a finite dimensional vector space over $E_{\Lambda}$ and $\pi:G\rightarrow{\rm GL}(V)$ is a continuous homomorphism.
@@
Let $dim_{E_{\Lambda}}(V)=m$. Now from [3, Proposition 9.3.5], there is a basis of $V$ such that $\pi(G)\subset{\rm GL}_{m}(\mathcal{O}_{\Lambda})$ for all $g\in G$. Let us consider the maps $\pi^{\prime}:G\rightarrow{\rm GL}_{m}(\mathbb{F}_{\Lambda})$ and $\tilde{\pi}:G\rightarrow{\rm PGL}_{m}(\mathbb{F}_{\Lambda})$ obtained from $\pi$ through the quotients maps $\mathcal{O}_{\Lambda}\rightarrow\mathbb{F}_{\Lambda}$ and ${\rm GL}_{m}(\mathbb{F}_{\Lambda})\rightarrow{\rm PGL}_{m}(\mathbb{F}_{\Lambda})$.
@@
[…]
@@
Let $\sigma=\pi|_{H}:H\rightarrow{\rm GL}(W)$. Then $\sigma^{\prime}:H\rightarrow{\rm GL}_{m}(\mathbb{F}_{\Lambda})$ and $\tilde{\sigma}:H\rightarrow{\rm PGL}_{m}(\mathbb{F}_{\Lambda})$ and ${\rm Image}(\tilde{\sigma})=\tilde{\pi}(H)$. Also, $\tilde{\pi}(G):\tilde{\pi}(H)]\leq[G:H]$.
The following discussion is valid for a general context of a finite dimensional representation of a finite index subgroup $H$ of a group $G$. In this manuscript, we only consider the case of ${\rm Gal}(\bar{\mathbb{Q}}/\mathbb{Q})$ and ${\rm Gal}(\bar{K}/K)$.
@@
[…]
@@
Let $H$ be a normal subgroup of $G$ (i.e., $K$ is Galois over $\mathbb{Q}$}) such that $[G:H]=n$. Since $dim(W)=m$, we have $dim(V)=mn$. Without loss of generality, let $s_{1}=1$. We label the basis of $V$ as
$\begin{split}f_{1}=\phi_{1,w_{1}},\ f_{2}=\phi_{1,w_{2}},\ \dots,\ f_{m}=\phi_{1,w_{m}},\\ f_{m+1}=\phi_{s_{2},w_{1}},\ \dots,\ f_{2m}=\phi_{s_{2},w_{m}},\dots,\\ f_{(n-1)m+1}=\phi_{s_{n},w_{1}},\ \dots,\ f_{nm}=\phi_{s_{n},w_{m}}.\end{split}$
@@
Since $\pi(G)\subset{\rm GL}_{m}(\mathcal{O}_{\Lambda})$, we conclude that $\rho(G)\subset{\rm GL}_{nm}(\mathcal{O}_{\lambda})$ and $\rho^{\prime}:G\rightarrow{\rm GL}_{nm}(\mathbb{F}_{\Lambda})$ and $\tilde{\rho}:G\rightarrow{\rm PGL}_{nm}(\mathbb{F}_{\Lambda})$ can be defined.
@@
All the following results for the pair ($\rho$, $\pi$) in this subsection also hold for pairs ($\rho^{\prime}$, $\pi^{\prime}$) as well as ($\tilde{\rho}$, $\tilde{\pi}$).
@@
Define a map from $\rho(H)$ to $\pi(H)$ by first block projection sending $\rho(h)$ to $\pi(h)$. ∎
@@
$1\rightarrow\rho(H)\rightarrow\rho(G)\rightarrow G/H\rightarrow 1.$
@@
Let $G/H$ be cyclic with representatives of $H$-cosets in $G$ of the form $\{1,s,s^{2},\dots,s^{n-1}\}$. If $g=hs^{i}\in H$ for $0\leq i\leq n-1$, the matrix $\rho(hs^{i})$ is given by
@@
(3.1) $1\rightarrow\rho(H)\rightarrow\rho(G)\rightarrow G/H\rightarrow 1$
is right split (that is $\rho(G)\cong\rho(H)\rtimes G/H$ for some semidirect product group law) if and only if there exists a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ such that $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(G)$. (In fact it is a subgroup with same group structure as $\{s_{i}H\}_{i}=G/H$)
@@
Let $\{r_{1},r_{2},...,r_{n}\}$ be a set of representatives of $H$-cosets in $G$. If the exact sequence is right split, let $\iota$ be splitting with $\gamma\circ\iota=id_{G/H}$. Hence $\gamma\circ\iota(r_{i}H)=r_{i}H$ for each $i$. Thus for each $i$, $\iota(r_{i}H)=\rho(h_{i}r_{i})$ for some $h_{i}\in H$. Let $s_{i}=h_{i}r_{i}$ for each $i$. Hence $\{s_{1},s_{2},...,s_{n}\}$ also forms a set of representatives for $H$-cosets in $G$ with $s_{i}H=r_{i}H$ and $\iota(s_{i}H)=\rho(s_{i})$ for each $i$. Since $\iota$ is a homomorphism, $\{\rho(s_{i})\}_{i}$ forms a multiplicatively closed subset of $\rho(G)$.
Conversely, suppose there exists a set of representatives of $H$-cosets in $G$, $\{s_{1},s_{2},\dots,s_{n}\}$ with $s_{1}\in H$ and $\{\rho(s_{i})\}_{i}$ forming a multiplicatively closed subset of $\rho(G)$. Hence $\{\rho(s_{i})\}_{i}$ forms a subgroup of $\rho(G)$. $\{\rho(s_{i})\}_{i}$ is a group with the same group structure as $\{s_{i}H\}_{i}=G/H$. Then we can define $\iota(s_{i}H)=\rho(s_{i})$. Hence $\iota$ is a homomorphism. Also, $\gamma(\iota(s_{i}H))=\gamma(\rho(s_{i}))=s_{i}H$ hence $\gamma\circ\iota=id_{G/H}$. Hence the exact sequence is right split. ∎
@@
We can similarly prove the statements in Remark 2.4.1 in previous section.
@@
If there exists a set of coset representatives of $H$ in $G$ which forms a multiplicatively closed set, then the exact sequence 3.1 is right split.
@@
Let $K$ be a cyclic extension of $\mathbb{Q}$ and let $G$ and $H$ be as above. Then the exact sequence 3.1 is right split if and only if there exists a set of representatives of $H$-cosets in $G$ of the form $\{1,s,s^{2},\dots,s^{n-1}\}$ with $\rho(s)^{n}=1$. Also, if these statements are true then $order(\rho(s))=n$ and $s^{n}\in H$.